Janice can mow the lawn in 2 hours. Her friend can mow the same lawn in 2 hours. How long would it take them to mow the lawn if they had two lawn mowers and worked for the same lenght of time?
Answer by stanbon(75887)

Janice can mow the lawn in 2 hours. Her friend can mow the same lawn in 2 hours. How long would it take them to mow the lawn if they had two lawn mowers and worked for the same lenght of time?
-----------------------------------
Janice DATA:
time = 2 hrs/job ; rate = 1/2 job/hr
----
Friend DATA:
time = 2 hrs/job ; rate = 1/2 job/hr
---
Together DATA:
time = x hrs/job ; rate = 1/x job/hr
-----
Equation:
rate + rate = together rate
1/2 + 1/2 = 1/x
1/x = 1
x = 1 hr (time for them to do the job together)
-------------------------------
Cheers,
Stan H.